Operations Report – February 9th

Crew 204 Operations Report 09 – FEB – 2019

SOL: 12

Name of person filing report: Kunal Naik (Space_bot)

Non-nominal systems: Generator

Notes on non-nominal systems: Generator: Non- nominal. The alternator needs to be replaced and the coolant system flushed and hoses replaced. All parts already arrived.

Generator (hours run): Turned on this evening by 1700 (9th Feb), Stopped by this morning 0830 (9th Feb)

Solar— SOC % – Turned off Gen (9th Feb) 72%, Turned on Gen (9th Feb) 55%

Diesel Reading – 28%

Propane Reading – ~ 50%

Ethanol Free Gasoline – 10 gallons

Water (Auxillary tank) – Not in use gallons

Water (Static tank) – around 50%, ~ 250 gallons

Auxillary to Static tank transfer – No

Gallons transferred: Not Applicable

Water in GreenHab – ~ 159 gallons

Water (loft) – 12 marked level

Static to Loft Pump used – yes

Water Meter: 01409959 units

Toilet tank emptied: yes

Deimos rover used: No, still not functional

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not Applicable

Beginning charge: Not Applicable

Ending charge: Not Applicable

Currently charging: Not Applicable

Opportunity rover used: Not Used

Hours: 48.5 hours

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Curiosity rover used: Not Used

Hours: 73.6 hours

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Spirit rover used: Not used

Hours: 72.8

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Notes on rovers: Curiosity returned yesterday with a very low charge, compared with Oppy. For long distance EVA., it’s recommended to use Oppy.

ATV’s Used: None

Reason for use: Not used

Oil Added? N/A

ATV Fuel Used: Nope

# Hours the ATVs were Used today: None

Notes on ATVs: None

HabCar used and why, where? To town for Dinner

Crew Car used and why, where? To town for Dinner

General notes and comments: HabCar: back on campus. The car was taken to the workshop in order to fix the transmission issue. It was successfully repaired and ready to work. Currently, it doesn’t have a spare tire. The tire is on its way. One of the rear tires had a nail on it, it’s already patched.

Summary of internet: Nothing to report

Summary of suits and radios: Nothing to report

Summary of Hab operations: Nothing to report.

Summary of Green Hab operations: Nothing to report

Summary of Science Dome operations: Nothing to report

Summary of RAMM operations: Nothing to report

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report.

Questions, concerns and requests to Mission Support: None